For a particular pickup truck the percent markup is known to be 115% based on cost to the seller. If the seller paid $15,800 for the truck, what would be the percent markup be based on the sale price? (Round to the nearest tenth percent
Answer by Theo(13342)

the markup is 115% based on the cost to the seller.

if the seller paid 15,800 for the truck, then the seller takes 115% of that and adds it to the price that he sells the truck for.

15,800 + 115% * 15,800 = 15,800 * 115 / 100 * 15,800 = 15,800 * 1.15 = 18,170.

15,800 is the cost and 18,170 is the markup.

the selling price of the truck is 15,800 + 18,170 = 33,970.